Deposits of 5.800$ are made every 6 months into a fund starting today and continuing until 10 deposits have been made. Six months after the last deposit there is 70.630$ in the fund. What nominal rate of interest, j2 do the deposits earn?
1
Expert's answer
2021-10-28T16:35:02-0400

Future value of the annual deposit

"= contribution \times (\frac {(1+r)^{p}-1}{r})"

r is rate of interest

p is no. of period

nominal rate of interest j2 = r*2

p=10

contribution=5800

final value =70630

"70630=5800(\frac {(1+r)^{10}-1}{r}) \\

 \frac{70630}{5800}=\frac {(1+r)^{10}-1}{r} \\

12.177 \times r = (1+r)^{10} -1 \\

r ≈ 0.04304"

nominal rate of interest j2 "= 0.04304 \times 2 = 0.08608 = 8.6 \; \%"